# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v2

@@ 3.1. Variant Isoperimetry
$A^{c_{n}\leq t}\supseteq\tilde{A}^{c_{n}\leq t}$ $=\bigcup_{\mathbf{x}\in\tilde{A}}\{\mathbf{x}\}^{c_{n}\leq t}$
$\supseteq\bigcup_{\mathbf{u}\in\hat{A}}\Big(\{\mathbf{u}\}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}}\times A_{\mathbf{u}}^{c_{n^{\prime\prime}}\leq 2^{p}n\delta^{\prime}}\Big)$
$\supseteq\bigcup_{\mathbf{u}\in\hat{A}}\Big(\{\mathbf{u}\}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}}\times B\Big)$
$=\hat{A}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}}\times B.$
@@
$\limsup_{n\to\infty}E_{n}(\alpha,\tau)$
(3.35) $\leq\limsup_{n\to\infty}E_{n^{\prime}}(\alpha^{\prime\prime}+\ln\nu_{X}(F),\tau^{\prime}|\nu_{X|F},\nu_{Y})$
(3.36) $\leq\psi(\alpha^{\prime\prime}+\ln\nu_{X}(F),\tau^{\prime}|\nu_{X|F},\nu_{Y})$
(3.37) $\leq\psi(\alpha^{\prime\prime},\tau^{\prime}|\nu_{X},\nu_{Y}).$
@@
Since $\lim_{\alpha^{\prime}\downarrow\alpha}\psi(\alpha^{\prime},\tau)\geq\psi(\alpha,\tau)$, it suffices to prove $\lim_{\alpha^{\prime}\downarrow\alpha}\psi(\alpha^{\prime},\tau)\leq\psi(\alpha,\tau)$. Denote $\mathsf{C}=\mathsf{W}_{p}^{p}$. Denote
@@
To prove that $\alpha\mapsto\psi(\alpha,\tau)$ is continuous for $\tau\geq 0$, it suffices to prove that $\psi(\alpha_{k},\tau)\to\psi(\alpha,\tau)$ for all sequences $\alpha_{k}$ such that $\alpha_{k}\downarrow\alpha$. Denote $\pi_{XW}^{(k)}$ as a distribution approximately attaining the supremization in $\psi(\alpha_{k},\tau)$ within a gap $1/k$. Denote $\pi_{Y|W}^{(k)}$ as a conditional distribution attaining $g(\pi_{XW}^{(k)},\tau)$, i.e., the infimization in $\psi(\alpha_{k},\tau)$. Since $\pi_{XW}^{(k)}$ is in $\{\pi_{XW}:D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha\}$, a compact set, by passing a subsequence, we can assume that $\pi_{XW}^{(k)}$ converges to some distribution $\pi_{XW}^{*}$. Denote $\pi_{Y|W}^{*}$ as a conditional distribution attaining $g(\pi_{XW}^{*},\tau-\delta)$. Since by considering the feasible solution $\pi_{X|W}$ for the minimization in $g(\pi_{XW},\tau)$, one can observe that $g(\pi_{XW},\tau)$ is bounded by the finite number $\alpha$. So, without loss of generality, we assume that $D(\pi_{Y|W=w}^{*}\|\nu)$ is finite for both $w=0$ and $1$.
@@
So, $g(\pi_{XW}^{*},\tau-\delta)\leq\psi(\alpha,\tau-\delta)$, which, combined with (3.38), implies that
$\limsup_{k\to\infty}\psi(\alpha_{k},\tau)\leq\psi(\alpha,\tau-\delta).$
Letting $\delta\downarrow 0$ and by the continuity of $\tau\mapsto\psi(\alpha,\tau)$, it holds that $\limsup_{k\to\infty}\psi(\alpha_{k},\tau)\leq\psi(\alpha,\tau)$. ∎
3.2. From Transport to Variant Isoperimetry
We now turn back to the Riemannian manifold setting. Applying Corollary 5, we establish bounds on isoperimery from nonlinear log-Sobolev inequalities. Define
@@ 3.2. From Transport to Variant Isoperimetry
Theorem 11 (From Nonlinear Log-Sobolev to Isoperimetry).
Let $\mathcal{M}$ be a Riemannian manifold equipped with a reference probability measure $\nu=e^{-V}\mathrm{vol}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),V\in C^{2}(\mathcal{M})$. Assume (1.8) holds. Assume that holds the inequality
@@
for some convex $\theta$. Assume $\Upsilon(0,\alpha)<\infty$ for some (equivalently for all) $0<\alpha<\infty$. Let $E_{n}$ and $\psi$ be defined in Theorem 9 with $\nu=\nu_{X}=\nu_{Y}$ and $p=2$. Then, for any $\alpha,\tau\geq 0$,
@@
Obviously, the assumption in (1.8) implies $\mathbb{E}_{\nu}[d^{q}(X,x_{0})]<\infty$ for any $q\geq 1$. Denote $\alpha_{i}=D(\pi_{X|W=i}\|\nu)$, $i\in\{0,1\}$. By Theorem 5, given $\pi_{X|W}$ satisfying $\pi_{X|W=i}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),i\in\{0,1\}$ and given every $s_{i}\leq\alpha_{i}$, there is $\pi_{Y|W}$ such that $\pi_{Y|W=i}\in P_{2}^{\mathrm{ac}}(\mathcal{M})$, $s_{i}=D(\pi_{Y|W=i}\|\nu)$, and
@@
We now prove that there is a unique solution to the equations (3.44) and (3.45) (or (3.41) and (3.42)). Observe that given any $\alpha_{i}$,
$s\mapsto\frac{\Upsilon(s,\alpha_{i})}{\sqrt{\theta(s)}}$
is decreasing in $s\in(0,\alpha_{i}]$. So, from (3.45), for each $s_{1}$, there is unique $s_{0}=s_{0}^{*}(s_{1})$ satisfying (3.45). Moreover, $s_{0}^{*}(s_{1})$ is increasing in $s_{1}$, which implies that
@@
is decreasing in $s_{1}\in(0,\alpha_{1}]$. So, there is a unique solution $s_{1}=s_{1}^{*}$ ensuring this function equal to $\tau$, which in turn implies that $(s_{1}^{*},s_{0}^{*}(s_{1}^{*})$) is the unique solution to the equations in (3.44) and (3.45). This yields (3.39) under the assumption $\lambda\in(0,1)$.
@@
If $\lambda=1$, then the optimal $s_{1}^{*}$ for the inner minimization is $s_{1}^{*}=\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha_{1})$, and $s_{0}^{*}$ can be arbitrary, where $\Upsilon_{1}^{-1}$ is the inverse of $\Upsilon$ w.r.t. the first argument when the second is fixed to be $\alpha_{1}$. Similarly, if $\lambda=0$, then the optimal $s_{0}^{*}$ for the inner minimization is $s_{0}^{*}=\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha_{0})$ and $s_{1}^{*}$ can be arbitrary. For both these two cases, the upper bound in (3.43) cannot be greater than $\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha)$. By continuity, one can find that the most right hand side in (3.39) is greater than or equal to $\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha)$. So, the inequality in (3.39) still holds for $\lambda=0$ or $1$. ∎
3.3. Interpretation from Concentration of Measure
The isoperimetric problem with thick boundary admits a natural interpretation from the perspective of concentration of measure. Concentration of measure principle states that for any set of not too small probability, slightly enlarging this set will always have probability close to one. Formally, let $r(a)$ be the minimum number such that the $r$-enlargement of any set of measure $a$ will always have probability close to one. Denote $a=e^{-n\alpha}$ and $r^{2}(a)=n\tau(\alpha)$. By our results, for sufficiently large $n$, $\tau(\alpha)$ is the minimum $\tau$ such that $\psi(\alpha,\tau)=0$, where $\mathsf{C}=\mathsf{W}_{2}^{2}$. Solving it yields
$\tau\geq\mathsf{C}(\pi_{X|W},\nu|\pi_{W})$
for all $\pi_{XW}$ such that $D(\pi_{X|W}\|\nu|\pi_{W})\leq\alpha$. Hence, $\tau(\alpha)$ satisfies
which is the upper concave envelope of $\tau_{0}$ given below
Recall that $K_{C}$ is the optimal transport constant in the following inequality:
$\mathsf{W}_{2}(\pi,\nu)\leq\sqrt{\frac{2}{K_{C}}D(\pi\|\nu)},\;\forall\pi.$
Hence,
$\tau(\alpha)\leq\frac{2\alpha}{K_{C}},\;\forall\alpha>0,$
and the constant $K_{C}$ cannot be improved. This implies that for sufficiently large $n$, enlarging any set of measure $e^{-\frac{K_{C}r^{2}}{2}}$ by the distance $r$ will always make it have probability close to one, and moreover, the constant $K_{C}$ cannot be improved. In contrast, another (but indeed equivalent) interpretation of $K_{C}$ in the concentration of measure is that for all $A$ such that $\nu_{n}(A)=1/2$, it always holds that
$\nu_{n}((A^{r})^{c})\leq e^{-\frac{K_{C}r^{2}}{2}},\;\forall r>0$
for some constant $C$. Here, the constant $K_{C}$ cannot be improved either.
4. From Variant Isoperimetry to Classic Isoperimetry
4.1. Connection between Variant Isoperimetry and Classic Isoperimetry
$\nu_{n}^{+}(A)$ $=\liminf_{r\downarrow 0}\frac{\nu_{n}(A^{r})-\nu_{n}(A)}{\ln\nu_{n}(A^{r})-\ln\nu_{n}(A)}$
(4.1) $\qquad\qquad\times\frac{\ln\nu_{n}(A^{r})-\ln\nu_{n}(A)}{r}$
However, to apply Theorem 11, we need keep $r$ to be large (linear in $n$), and hence, we need remove $\liminf_{r\downarrow 0}$ in the last line. To this end, we need some nice properties of isoperimetric minimizers from geometric measure theory.
Definition 4.
Bayle [6] and Morgan [34, 36] showed that isoperimetric minimizers always exist and the isoperimetric profile is concave. See an explicit explanation in [32, p. 216].
Theorem 12 (Bayle [6], Morgan [34, 36]).
Assume that (1.5) holds. Then, the following hold.
  1. (1)
    For any $a\in(0,1)$, there always exists an open isoperimetric minimizer $A$ in $(\mathcal{M}^{n},\nu_{n})$ of given measure $a$ such that: (a) its boundary $\partial A$ consists of disjoint regular part and singular part; (b) the regular part is a smooth hypersurface which has a constant $\nu_{n}$-total curvature, denoted by $H_{\nu_{n}}(A)$; (c) the singular part has dimension strictly lower than that of the regular part, more precisely, lower than or equal to the dimension of the regular part minus $7$.
  2. (2)
    The isoperimetric profile $I_{n}$ is a concave function on $[0,1]$.
  3. (3)
Theorem 13 (Morgan [34, Theorem 2, Remark 3]).
Assume that for $\kappa\in\mathbb{R}$,
$\mathrm{Ric}_{g}+\mathrm{Hess}_{g}V\geq\kappa g.$
Note that the sign of the total curvature $H_{S}^{\sigma}(p)$ here is opposite to the one in [34, Theorem 2, Remark 3].
Based on the two theorems above, we build a bridge between the classic isoperimetric problem and the isoperimetric problem with thick boundary.
Theorem 14.
Assume that (1.5) holds. Let $A\subseteq\mathcal{M}^{n}$ denote an isoperimetric minimizer in $(\mathcal{M}^{n},\nu_{n})$ of given measure $a\in(0,1)$ given in Theorem 12. Then, for any $r>0$,
@@ 4.1. Connection between Variant Isoperimetry and Classic Isoperimetry
$P(r)$ $:=\nu_{n}^{+}(A)\exp\left(H_{\nu_{n}}(A)r\right),$
$V(r)$ $:=\nu_{n}(A)+\int_{0}^{r}P(t){\text{\rm d}}t.$
Define
g(r):=V(r)P(0)P(r)V(0),g(r):=V(r)P(0)-P(r)V(0),
whose derivative satisfies
$g^{\prime}(r)$ $=P(r)P(0)-P^{\prime}(r)V(0)$
(4.3) $=P(r)P(0)-H_{\nu_{n}}(A)P(r)V(0).$
From Theorem 12, it is known that
where $a=\nu_{n}(A)$. Since $I_{n}$ is concave,
Hence,
$H_{\nu_{n}}(A)\leq\frac{P(0)}{V(0)},$
which, combined with (4.3), implies
$g^{\prime}(r)\geq 0.$
Observe that $g(0)=0$. So, for all $r\geq 0$, $g(r)\geq 0,$ i.e.,
$\frac{P(r)}{V(r)}\leq\frac{P(0)}{V(0)}.$
Noting that $P(r)=V^{\prime}(r)$, we then have that
$\ln\frac{V(r)}{V(0)}\leq\frac{P(0)}{V(0)}r.$
From Theorem 13, it is known that $V(r)\geq\nu_{n}(A^{r})$. We hence conclude that
4.2. Lower Bounds in Large Deviations Theorem
Define
where $\psi$ is defined in (3.17) but with $\nu=\nu_{X}=\nu_{Y}$ and $p=2$. As a consequence of Theorems 9 and 14, as well as Corollary 5, we can lower bound the asymptotics of the isoperimetric profile $I_{n}$ in terms of $\xi$ or $\breve{\Theta}$. Recall that $\breve{\Theta}$ defined in (1.7) is the function characterizing the optimal tradeoff in the nonlinear log-Sobolev inequality. We obtain the following theorem which is the lower bound part of Theorem 1.
Theorem 15 (From Nonlinear Log-Sobolev to Isoperimetry).
Assume that (1.5) and (1.8) hold. Then, for any $\alpha>0$,
(4.5) $\liminf_{n\to\infty}\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}\sqrt{n}}$ $\geq\xi(\alpha)\geq\sqrt{\breve{\Theta}(\alpha)}.$
Proof.
Let $A$ be an isoperimetric minimizer of size $e^{-n\alpha}$ in the $n$-dimensional space. From Theorem 14, we see that for any $r>0$,
$\frac{\ln[\nu_{n}(A^{r})/\nu_{n}(A)]}{r}\leq\frac{\nu_{n}^{+}(A)}{\nu_{n}(A)}=\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}}.$
@@ 4.2. Lower Bounds in Large Deviations Theorem
[…]
$=\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}},$
where the last line follows from Theorem 9. Letting $\tau\to 0$, we obtain the first inequality in (4.5).
We now prove the second inequality in (4.5). If $\liminf_{n\to\infty}\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}\sqrt{n}}=\infty$, then there is nothing to prove. So, we assume $\liminf_{n\to\infty}\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}\sqrt{n}}<\infty$. By the first inequality in (4.5), $\liminf_{\tau\downarrow 0}\psi(\alpha,\tau)=\alpha$. Since $\psi(\alpha,\tau)\leq\alpha$, it holds that $\lim_{\tau\downarrow 0}\psi(\alpha,\tau)=\alpha$. On the other hand, by Corollary 5,
@@
$\xi(\alpha)$ $\geq\inf_{\begin{subarray}{c}\lambda\in(0,1),\alpha_{0},\alpha_{1}\geq 0:\\ \lambda\alpha_{1}+(1-\lambda)\alpha_{0}=\alpha,\\ \lambda s_{1}^{*}+(1-\lambda)s_{0}^{*}\geq\alpha-\epsilon\end{subarray}}\frac{\alpha-\left(\lambda s_{1}^{*}+(1-\lambda)s_{0}^{*}\right)}{\sqrt{\tau}}$
$=\inf_{\begin{subarray}{c}\lambda\in(0,1),\alpha_{0},\alpha_{1}\geq 0:\\ \lambda\alpha_{1}+(1-\lambda)\alpha_{0}=\alpha,\\ \lambda s_{1}^{*}+(1-\lambda)s_{0}^{*}\geq\alpha-\epsilon\end{subarray}}\frac{\lambda\left(\alpha_{1}-s_{1}^{*}\right)+(1-\lambda)\left(\alpha_{0}-s_{0}^{*}\right)}{\sqrt{\tau}},$
@@
5. From Classic Isoperimetry to Log-Sobolev
We now prove the other direction (i.e., the upper bound) in Theorem 1. Recall $\xi$ defined in (4.4) and $\bar{\Lambda}$ defined in (1.4).

### v3

@@ 3.1. From Transport to Variant Isoperimetry
$A^{c_{n}\leq t}\supseteq\tilde{A}^{c_{n}\leq t}$ $=\bigcup_{\mathbf{x}\in\tilde{A}}\{\mathbf{x}\}^{c_{n}\leq t}\supseteq\bigcup_{\mathbf{u}\in\hat{A}}\Big(\{\mathbf{u}\}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}}\times A_{\mathbf{u}}^{c_{n^{\prime\prime}}\leq 2^{p}n\delta^{\prime}}\Big)$
$\supseteq\bigcup_{\mathbf{u}\in\hat{A}}\Big(\{\mathbf{u}\}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}}\times B\Big)=\hat{A}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}}\times B.$
@@
(3.34) $\limsup_{n\to\infty}E_{n}(\alpha,\tau)$ $\leq\limsup_{n\to\infty}E_{n^{\prime}}(\alpha^{\prime\prime}+\ln\nu_{X}(F),\tau^{\prime}|\nu_{X|F},\nu_{Y})$
$\leq\psi(\alpha^{\prime\prime}+\ln\nu_{X}(F),\tau^{\prime}|\nu_{X|F},\nu_{Y})\leq\psi(\alpha^{\prime\prime},\tau^{\prime}|\nu_{X},\nu_{Y}).$
@@
Since $\lim_{\alpha^{\prime}\downarrow\alpha}\psi(\alpha^{\prime},\tau)\geq\psi(\alpha,\tau)$, it suffices to prove $\lim_{\alpha^{\prime}\downarrow\alpha}\psi(\alpha^{\prime},\tau)\leq\psi(\alpha,\tau)$. We now use the standard lower semi-continuity and compactness arguments to prove this point.
Denote $\mathsf{C}=\mathsf{W}_{p}^{p}$. Denote
@@
To prove that $\alpha\mapsto\psi(\alpha,\tau)$ is continuous for $\tau\geq 0$, it suffices to prove that $\psi(\alpha_{k},\tau)\to\psi(\alpha,\tau)$ for all sequences $\alpha_{k}$ such that $\alpha_{k}\downarrow\alpha$. Denote $\pi_{XW}^{(k)}$ as a distribution approximately attaining the supremization in $\psi(\alpha_{k},\tau)$ within a gap $1/k$. Denote $\pi_{Y|W}^{(k)}$ as a conditional distribution attaining $g(\pi_{XW}^{(k)},\tau)$, i.e., the infimization in $\psi(\alpha_{k},\tau)$. Since $\pi_{XW}^{(k)}$ is in $\{\pi_{XW}:D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha\}$, a weakly compact set (e.g., [25, Lemma 6.2.12]), by passing a subsequence, we can assume that $\pi_{XW}^{(k)}$ converges to some distribution $\pi_{XW}^{*}$. Denote $\pi_{Y|W}^{*}$ as a conditional distribution attaining $g(\pi_{XW}^{*},\tau-\delta)$. Since by considering the feasible solution $\pi_{X|W}$ for the minimization in $g(\pi_{XW},\tau)$, one can observe that $g(\pi_{XW},\tau)$ is bounded by the finite number $\alpha$. So, without loss of generality, we assume that $D(\pi_{Y|W=w}^{*}\|\nu)$ is finite for both $w=0$ and $1$.
@@
So, $g(\pi_{XW}^{*},\tau-\delta)\leq\psi(\alpha,\tau-\delta)$, which, combined with (3.35), implies that $\limsup_{k\to\infty}\psi(\alpha_{k},\tau)\leq\psi(\alpha,\tau-\delta).$ Letting $\delta\downarrow 0$ and by the continuity of $\tau\mapsto\psi(\alpha,\tau)$, it holds that $\limsup_{k\to\infty}\psi(\alpha_{k},\tau)\leq\psi(\alpha,\tau)$. ∎
3.2. Interpretation from Concentration of Measure
The isoperimetric problem with thick boundary admits a natural interpretation from the perspective of concentration of measure. Concentration of measure principle states that for any set of not too small probability, slightly enlarging this set will always have probability close to one. Formally, let $r(a)$ be the minimum number such that the $r$-enlargement of any set of measure $a$ will always have probability close to one. Denote $a=e^{-n\alpha}$ and $r^{2}(a)=n\tau(\alpha)$. By our results, for sufficiently large $n$, $\tau(\alpha)$ is the minimum $\tau$ such that $\psi(\alpha,\tau)=0$, where $\mathsf{C}=\mathsf{W}_{2}^{2}$. Solving it yields $\tau\geq\mathsf{C}(\pi_{X|W},\nu|\pi_{W})$ for all $\pi_{XW}$ such that $D(\pi_{X|W}\|\nu|\pi_{W})\leq\alpha$. Hence, $\tau(\alpha)$ satisfies
which is the upper concave envelope of $\tau_{0}$ given below
Recall that $K_{C}$ is the optimal transport constant in the following inequality: $\mathsf{W}_{2}(\pi,\nu)\leq\sqrt{\frac{2}{K_{C}}D(\pi\|\nu)},\;\forall\pi.$ Hence, $\tau(\alpha)\leq\frac{2\alpha}{K_{C}},\;\forall\alpha>0,$ and the constant $K_{C}$ cannot be improved. This implies that for sufficiently large $n$, enlarging any set of measure $\exp(-\frac{K_{C}r^{2}}{2})$ by the distance $r$ will always make it have probability close to one, and moreover, the constant $K_{C}$ cannot be improved. In contrast, another (but indeed equivalent) interpretation of $K_{C}$ in the concentration of measure is that for all $A$ such that $\nu_{n}(A)=1/2$, it always holds that $\nu_{n}((A^{r})^{c})\leq\exp(-\frac{K_{C}r^{2}}{2}),\;\forall r>0$ for some constant $C$. Here, the constant $K_{C}$ cannot be improved either.
4. From Variant Isoperimetry to Standard Isoperimetry
$\nu_{n}^{+}(A)$ $=\liminf_{r\downarrow 0}\frac{\nu_{n}(A^{r})-\nu_{n}(A)}{\ln\nu_{n}(A^{r})-\ln\nu_{n}(A)}\cdot\frac{\ln\nu_{n}(A^{r})-\ln\nu_{n}(A)}{r}$
However, to apply Theorem 14, we need keep $r$ to be large (in order of $\sqrt{n}$), and hence, we need remove $\liminf_{r\downarrow 0}$ in the last line. To this end, we need some nice properties of isoperimetric minimizers from geometric measure theory.
Definition 4.
Bayle [10] and Morgan [49, 51] showed that isoperimetric minimizers always exist and the isoperimetric profile is concave. See an explicit explanation in [45, p. 216].
Theorem 11 (Morgan [48, 49, 51] and Bayle [10]).
Under Assumption 2, the following hold.
  1. (1)
    For any $a\in(0,1)$, there always exists an open isoperimetric minimizer $A$ in $(\mathcal{M}^{n},\nu_{n})$ of given measure $a$ such that: (a) its boundary $\partial A$ consists of disjoint regular part and singular part; (b) the regular part is a $C^{2}$-smooth hypersurface (since $V\in C^{2}(\mathcal{M})$) which has a constant $\nu_{n}$-total curvature, denoted by $H_{\nu_{n}}(A)$; (c) the singular part has dimension strictly lower than that of the regular part, more precisely, lower than or equal to the dimension of the regular part minus $7$.
  2. (2)
    The isoperimetric profile $I_{n}$ is a concave function on $[0,1]$.
  3. (3)
Theorem 12 (Generalized Heintze-Karcher due to Morgan [49, Theorem 2, Remark 3]).
Assume that for $\kappa\in\mathbb{R}$,
$\mathrm{Ric}_{g}+\mathrm{Hess}_{g}V\geq\kappa g.$
Note that the sign of the total curvature $H_{S}^{\sigma}(p)$ here is opposite to the one in [49, Theorem 2, Remark 3]. Based on the two theorems above, we build a bridge between the standard isoperimetric problem and the isoperimetric problem with thick boundary.
Theorem 13.
Assume that Assumption 1 holds. Let $A\subseteq\mathcal{M}^{n}$ denote an isoperimetric minimizer in $(\mathcal{M}^{n},\nu_{n})$ of given measure $a\in(0,1)$ given in Theorem 11. Then, for any $r>0$,
Proof.
$P(r)$ $:=\nu_{n}^{+}(A)\exp\left(H_{\nu_{n}}(A)r\right),$
$V(r)$ $:=\nu_{n}(A)+\int_{0}^{r}P(t){\text{\rm d}}t.$
Define g(r):=V(r)P(0)P(r)V(0),g(r):=V(r)P(0)-P(r)V(0), whose derivative satisfies
(4.2) $g^{\prime}(r)=P(r)P(0)-P^{\prime}(r)V(0)=P(r)P(0)-H_{\nu_{n}}(A)P(r)V(0).$
From Theorem 11, it is known that
where $a=\nu_{n}(A)$. Since $I_{n}$ is concave,
Hence, $H_{\nu_{n}}(A)\leq\frac{P(0)}{V(0)},$ which, combined with (4.2), implies $g^{\prime}(r)\geq 0.$
Observe that $g(0)=0$. So, for all $r\geq 0$, it holds that $g(r)\geq 0,$ i.e., $\frac{P(r)}{V(r)}\leq\frac{P(0)}{V(0)}.$ Noting that $P(r)=V^{\prime}(r)$, we then have that $\ln\frac{V(r)}{V(0)}\leq\frac{P(0)}{V(0)}r.$ From Theorem 12, it is known that $V(r)\geq\nu_{n}(A^{r})$. We hence conclude that
5. Putting Everything Together: From Log-Sobolev to Standard Isoperimetry
In this section, we combine all the tools built in Sections 24 to prove the lower bound in our large deviations theorem, i.e., to deduce standard isoperimery from nonlinear log-Sobolev inequalities. To this end, we first deduce variant isoperimery from nonlinear log-Sobolev inequalities by combining Theorem 6 and Theorem 10. Define
@@ 5. Putting Everything Together: From Log-Sobolev to Standard Isoperimetry
Theorem 14 (From Nonlinear Log-Sobolev to Variant Isoperimetry).
Let $\mathcal{M}$ be a Riemannian manifold equipped with a reference probability measure $\nu=e^{-V}\mathrm{vol}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),V\in C^{2}(\mathcal{M})$. Assume that nonlinear log-Sobolev inequality
@@
holds for some convex $\theta$. Assume $\Upsilon(0,\alpha)<\infty$ for some (thus all) $0<\alpha<\infty$. Let $E_{n}$ and $\psi$ be defined in Theorem 10 (for variant isoperimetry) with $\nu=\nu_{X}=\nu_{Y}$ and $p=2$. Then, under Assumption 2, for any $\alpha,\tau\geq 0$,
@@
Obviously, Assumption 2 implies $\mathbb{E}_{\nu}[d^{q}(X,x_{0})]<\infty$ for any $q\geq 1$. By Theorem 10, the first inequality in (5.1) is obtained. We next prove the second inequality.
Denote $\alpha_{i}=D(\pi_{X|W=i}\|\nu)$, $i\in\{0,1\}$. By Theorem 6, given $\pi_{X|W}$ satisfying $\pi_{X|W=i}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),i\in\{0,1\}$ and given every $s_{i}\leq\alpha_{i}$, there is $\pi_{Y|W}$ such that $\pi_{Y|W=i}\in P_{2}^{\mathrm{ac}}(\mathcal{M})$, $s_{i}=D(\pi_{Y|W=i}\|\nu)$, and
@@
We now prove that there is a unique solution to the equations (5.6) and (5.7) (or (5.3) and (5.4)). Observe that given any $\alpha_{i}$, $s\mapsto\frac{\Upsilon(s,\alpha_{i})}{\sqrt{\theta(s)}}$ is decreasing in $s\in(0,\alpha_{i}]$. So, from (5.7), for each $s_{1}$, there is unique $s_{0}=s_{0}^{*}(s_{1})$ satisfying (5.7). Moreover, $s_{0}^{*}(s_{1})$ is increasing in $s_{1}$, which implies that
@@
is decreasing in $s_{1}\in(0,\alpha_{1}]$. So, there is a unique solution $s_{1}=s_{1}^{*}$ ensuring this function equal to $\tau$, which in turn implies that $(s_{1}^{*},s_{0}^{*}(s_{1}^{*})$) is the unique solution to the equations in (5.6) and (5.7). This yields the second inequality in (5.1) under the assumption $\lambda\in(0,1)$.
@@
If $\lambda=1$, then the optimal $s_{1}^{*}$ for the inner minimization is $s_{1}^{*}=\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha_{1})$, and $s_{0}^{*}$ can be arbitrary, where $\Upsilon_{1}^{-1}$ is the inverse of $\Upsilon$ w.r.t. the first argument when the second is fixed to be $\alpha_{1}$. Similarly, if $\lambda=0$, then the optimal $s_{0}^{*}$ for the inner minimization is $s_{0}^{*}=\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha_{0})$ and $s_{1}^{*}$ can be arbitrary. For both these two cases, the upper bound in (5.5) cannot be greater than $\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha)$. By continuity, one can find that the most right hand side in (5.1) is greater than or equal to $\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha)$. So, the second inequality in (5.1) still holds for $\lambda=0$ or $1$. ∎
We now deduce standard isoperimery from nonlinear log-Sobolev inequalities. Define
where $\psi$ is defined in (3.17) but with $\nu=\nu_{X}=\nu_{Y}$ and $p=2$. As a consequence of Theorem 13 and Theorem 10 (or Theorem 14), we can lower bound the asymptotics of the isoperimetric profile $I_{n}$ in terms of $\xi$ (or $\breve{\Theta}$). Recall that $\breve{\Theta}$ defined in (1.7) is the function characterizing the optimal tradeoff in the nonlinear log-Sobolev inequality. We obtain the lower bound part of Theorem 1. Recall $\underline{\Lambda}$ defined in (1.5).
Theorem 15 (From Nonlinear Log-Sobolev to Standard Isoperimetry).
Under Assumptions 1 and 2, for any $\alpha>0$,
(5.9) $\underline{\Lambda}(\alpha)$ $\geq\xi(\alpha)\geq\sqrt{\breve{\Theta}(\alpha)}.$
@@
Let $A_{n}$ be an isoperimetric minimizer of size $e^{-n\alpha}$ in the $n$-dimensional space. From Theorem 13, we see that for any $r>0$,
$\frac{\ln[\nu_{n}(A_{n}^{r})/\nu_{n}(A_{n})]}{r}\leq\frac{\nu_{n}^{+}(A_{n})}{\nu_{n}(A_{n})}=\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}}.$
@@
$\underline{\Lambda}(\alpha)$ $\geq\liminf_{n\to\infty}\frac{\ln[\nu_{n}(A_{n}^{\sqrt{n\tau}})/\nu_{n}(A_{n})]}{n\sqrt{\tau}}\geq\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}},$
where the last inequality follows from Theorem 10. Letting $\tau\to 0$, we obtain the first inequality in (5.9).
We now prove the second inequality in (5.9). If $\underline{\Lambda}(\alpha)=\infty$, then there is nothing to prove. So, we assume $\underline{\Lambda}(\alpha)<\infty$. By the first inequality in (5.9), $\liminf_{\tau\downarrow 0}\psi(\alpha,\tau)=\alpha$. Since $\psi(\alpha,\tau)\leq\alpha$, it holds that $\lim_{\tau\downarrow 0}\psi(\alpha,\tau)=\alpha$. On the other hand, by Theorem 14,
@@
$\xi(\alpha)$ $\geq\inf_{\begin{subarray}{c}\lambda\in(0,1),\alpha_{0},\alpha_{1}\geq 0:\\ \lambda\alpha_{1}+(1-\lambda)\alpha_{0}=\alpha,\\ \lambda s_{1}^{*}+(1-\lambda)s_{0}^{*}\geq\alpha-\epsilon\end{subarray}}\frac{\lambda\left(\alpha_{1}-s_{1}^{*}\right)+(1-\lambda)\left(\alpha_{0}-s_{0}^{*}\right)}{\sqrt{\tau}},$
@@
6. The Other Direction: From Standard Isoperimetry to Log-Sobolev
We now prove the other direction (i.e., the upper bound) in Theorem 1. Recall $\xi$ defined in (5.8) and $\bar{\Lambda}$ defined around (1.5).